# Bridging Quantum Computing and Nuclear Structure: Atomic Nuclei on a Trapped-Ion Quantum Computer

### v1

@@ Methods
One can reproduce the results of this work (within uncertainties due to randomness) by the code available at the GitHub repository [58]. For classical computations, the data can be reproduced by using the NuclearToolkit.jl package [40, 41] with the interaction files available through KSHELL [38, 39]. We utilized the Qiskit [59], PennyLane [55], and pytket [53, 54] for quantum computations. Experiments on the Reimei device were carried out using the Quantinuum’s Nexus [28] platform.
Acknowledgements.
@@
This research was supported in part by JSPS Grant-in-Aid for Scientific Research (Grant Nos. JP22K14030, JP25H01511, JP25K01688), JST COI-NEXT (Grant No. JPMJPF2221), MEXT Q-LEAP (Grant No. JPMXS0118067246), and RIKEN TRIP initiative (Nuclear Transmutation). This paper is based on results obtained from a project, JPNP20017, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
Author contributions
S.Y. conceived the idea of the project, developed the framework and the code, and performed the simulations and analyses. T.S. and T.O. contributed to the design of the ansatz and of the quantum circuits from chemical background. M.K. supervised the project. All authors contributed to verifying the results and writing the manuscript.
Supplementary Information
I Derivation of hard-core-boson form of the Hamiltonian
@@
Here we provide the detailed derivation of the pair-wise Hamiltonians for the systems of interest, corresponding to Eq. (3) of the main text. Starting from the fermionic Hamiltonian with up to two-body interactions, we introduced time-reversal pairs $(i,\bar{i})$ and defined pair creation, annihilation, and number operators in Eqs. (5)–(6).
@@
$\frac{1}{4}\sum_{i,j\notin\{i,\bar{i}\}}V_{ijij}a^{\dagger}_{i}a^{\dagger}_{j}a_{j}a_{i}=\frac{1}{4}\sum_{i,j\notin\{i,\bar{i}\}}V_{ijij}n_{i}n_{j}=\sum_{i<0}\sum_{j\notin\{i,\bar{i}\}<0}(V_{ijij}+V_{i\bar{j}i\bar{j}})n_{i}n_{j}=\frac{1}{2}\sum_{\tilde{i}\tilde{j}}V^{m}_{\tilde{i}\tilde{j}}n_{i}n_{j},$ (22)
where the indices $i,j,k,l$ denote a Fermionic single-particle states having $\{n,l,j,j_{z},t_{z}\}$ quanta. The $\tilde{i},\tilde{j}$ represent pair-wise basis states obtained by folding single-particle states $i$ and $\bar{i}$ (the time-reversed partner).
II Uncertainty Bands in Figure 1
Figure 6: Uncertainty estimation for pUCCD (G) results. Sampled possible realization of ansatz is fitted by a Gamma distribution for the pUCCD results of a 20O, b 50Ca and c 66Ni.
@@ II Uncertainty Bands in Figure 1
The fitted Gamma distributions were then used to construct the shaded uncertainty bands in Fig. 1 of the main text. We note that the pUCCD (GS) ansatz typically yields results close to the optimal edge of the pUCCD (G) distribution, except for cases such as 24O, 50Ca, and 52Ca where the qubit ordering or single-particle energy structure leads to less favorable initialization.
III Results of the pUCCD (GS) ansatz on IBM’s noisy simulator
@@
For completeness, we benchmark the pUCCD (GS) ansatz on IBM’s noisy simulator (FakeTorino backend), which reflects the heavy-hexagon connectivity of IBM devices. In this setting, additional SWAP gates are required, increasing the CNOT count to $3N_{G}$ (where $N_{G}$ is the number of Givens rotations), compared to $2N_{G}$ under all-to-all connectivity. Figure 7 summarizes the relative errors with respect to the ideal statevector simulations. Three measurement strategies were compared:
  • Estimator method (blue triangles): direct expectation-value evaluation.

### v2

@@ IV Summary and outlook
The data presented in this manuscript, including circuits and encoded Hamiltonians, are available from the GitHub repository and the Zenodo archive [55]. For classical calculations, the exact diagonalization results can be reproduced by using the NuclearToolkit.jl package [26, *Repo_NuclearToolkit.jl] with the interaction files available through KSHELL [24, *KSHELL2]. We utilized the Qiskit [56], PennyLane [39], and pytket [44, *pytket-quantinuum] for quantum computations. Experiments on the Reimei device were carried out using the Quantinuum’s Nexus [46] platform.
acknowledgments
@@
This research was supported in part by JSPS Grant-in-Aid for Scientific Research (Grant Nos. JP22K14030, JP25H01511, JP25K01688), JST PRESTO (Grant No. JPMJPR25F8), JST COI-NEXT (Grant No. JPMJPF2221), MEXT Q-LEAP (Grant No. JPMXS0118067246), and RIKEN TRIP initiative (Nuclear Transmutation). This work is based on the results obtained from a project, JPNP20017, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
Supplemental Material
In this Supplemental Material, we provide supplementaray information regarding (i) the derivation of the hard-core-boson form of the nuclear Hamiltonian, (ii) the uncertainty bands of the pUCCD(G) ansatz, (iii) the results of the pUCCD(GS) ansatz on IBM’s noisy simulator, and (iv) description of Zenodo archive containing data analyzed in this work.
V Derivation of hard-core-boson form of the Hamiltonian
@@
Here we present the detailed derivation of the pair-wise Hamiltonians for the systems of interest, corresponding to Eq. (2) of the main text. Starting from the fermionic Hamiltonian with up to two-body interactions, we introduced time-reversal pairs $(i,\bar{i})$ and defined pair creation, annihilation, and number operators in Eqs. (4)–(5).
@@
$\frac{1}{4}\sum_{i,j\notin\{i,\bar{i}\}}V_{ijij}a^{\dagger}_{i}a^{\dagger}_{j}a_{j}a_{i}=\frac{1}{4}\sum_{i,j\notin\{i,\bar{i}\}}V_{ijij}n_{i}n_{j}$ (22)
where the indices $i,j,k,l$ denote a Fermionic single-particle states having $\{n,l,j,j_{z},t_{z}\}$ quanta. The $\tilde{i},\tilde{j}$ represent pair-wise basis states obtained by folding single-particle states $i$ and $\bar{i}$ (the time-reversed partner). Rewriting these in terms of pair operators defined in Eqs. (3)–(5) of the main text, we obtain
$\sum_{i<0}V_{i\bar{i}i\bar{i}}a^{\dagger}_{i}a^{\dagger}_{\bar{i}}a_{\bar{i}}a_{i}$ $\rightarrow\sum_{\tilde{i}}V^{p}_{\tilde{i}\tilde{i}}A^{\dagger}_{\tilde{i}}A_{\tilde{i}}$, (23)
$\sum_{i\neq j<0}V_{i\bar{i}j\bar{j}}a^{\dagger}_{i}a^{\dagger}_{\bar{i}}a_{\bar{j}}a_{j}$ $\rightarrow\sum_{\tilde{i}\neq\tilde{j}}V^{p}_{\tilde{i}\tilde{j}}A^{\dagger}_{\tilde{i}}A_{\tilde{j}}$, (24)
$\frac{1}{4}\sum_{i,j\notin\{i,\bar{i}\}}V_{ijij}n_{i}n_{j}$ $\rightarrow\sum_{\tilde{i}\neq\tilde{j}}V^{m}_{\tilde{i}\tilde{j}}N_{\tilde{i}}N_{\tilde{j}}$, (25)
where we put the superscripts $p$ and $m$ to distinguish pairing and monopole terms, respectively.
VI Uncertainty Bands of the pUCCD(G) ansatz
Figure 6: Uncertainty estimation for pUCCD(G) results. Sampled possible realization of ansatz is fitted by a Gamma distribution for the pUCCD results of (a) 20O, (b) 50Ca and (c) 66Ni.
@@ VI Uncertainty Bands of the pUCCD(G) ansatz
The fitted Gamma distributions were then used to construct the shaded uncertainty bands in Fig. (3) of the main text. We note that the pUCCD(GS) ansatz typically yields results close to the optimal edge of the pUCCD(G) distribution, except for cases such as 24O, 50Ca, and 52Ca where the qubit ordering or single-particle energy structure leads to less favorable initialization.
VII Results of the pUCCD(GS) ansatz on IBM’s noisy simulator
@@
For completeness, we benchmark the pUCCD(GS) ansatz on IBM’s noisy simulator (FakeTorino backend), which reflects the heavy-hexagon connectivity of IBM devices. In this setting, additional SWAP gates are required, increasing the number of two-qubit gates to $3N_{G}$ (where $N_{G}$ is the number of Givens rotations), compared to $2N_{G}$ under all-to-all connectivity. Figure 7 summarizes the relative errors with respect to the ideal statevector simulations. Three measurement strategies were compared:
  • Estimator (blue triangles): direct expectation-value evaluation.
@@
VIII Description of data archive
The data that support the findings of this study are openly available in:
One can find the following data in the repository and archive:
  • circuit_parameters: Optimized parameters for pUCCD ansatze.
  • Data_measured: measured data on IBM’s FakeTorino simulator and random sampling data for uncertainty estimation.
  • interaction_file: interaction files used in this work.
  • pytket_circuits: pytket circuit files for pUCCD ansatze.
  • qiskit_Hamiltonians: encoded Hamiltonians in qiskit format.
  • retrieved_data: measurement results retrieved from the Reimei quantum computer for both emulator and real hardware runs.